A stone is dropped from the top of a cliff and one second later a second stone is thrown vertically upward with a velocity of 20m/s. how far below the top of the cliff will the second stone overtake the first?


1
Expert's answer
2021-03-08T08:18:29-0500

Let's take the upwards as the positive direction. Then, we can write the vertical displacement of the first stone dropped from the top of the cliff:


"y_1=-\dfrac{1}{2}gt^2."

Similarly, we can write the vertical displacement of the second stone thrown vertically upward one second later:


"y_2=v_0(t-1)-\dfrac{1}{2}g(t-1)^2."

At the moment when the second stone overtake the first "y_1=y_2" and we can write:


"-\dfrac{1}{2}gt^2=v_0(t-1)-\dfrac{1}{2}g(t-1)^2,""-\dfrac{1}{2}gt^2=v_0t-v_0-\dfrac{1}{2}g(t^2-2t+1),""-\dfrac{1}{2}gt^2=v_0t-v_0-\dfrac{1}{2}gt^2+gt-\dfrac{1}{2}g,""20t-20+9.8t-4.9=0,""t=\dfrac{24.9}{29.8}=0.83\ s."

Finally, we can find how far below the top of the cliff will the second stone overtake the first:


"y_1=-\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(0.83\ s)^2=-3.4\ m."

The second stone overtake the first 3.4 m below the top of the cliff.